The speed of a plane is 120 miles per hour in a calm. With the wind it can cover a certain distance in 4 hours, but against the wind it can cover only 3/5 of that distance in the same time. Find the velocity of the wind.
Answer by nerdybill(7384)

The speed of a plane is 120 miles per hour in a calm. With the wind it can cover a certain distance in 4 hours, but against the wind it can cover only 3/5 of that distance in the same time. Find the velocity of the wind.
.
Let x=velocity of the wind
.
distance with wind = 4(120+x)
distance against the wind = 4(120-x)
.
(3/5)(4)(120+x) = 4(120-x)
(3/5)(120+x) = (120-x)
(3)(120+x) = 5(120-x)
360+3x = 600-5x
360+8x = 600
8x = 240
x = 30 mph